UT 20140201 - 20140228 LUCI2 Closed-dome Commissioning

Observer: LUCI-team
T elescope Operator: various
LBTO Support:

Details

February 4 2014

Test by JH

Dave reported problems with GEIRS when starting his Dark_RN_NoCU.xml script. Tried to check this.

=> Started GEIRS, readout image, no problem
=> Started readout manager: readout image => error message
=> Stopped: GEIRS readout, readout manager, scheduler and journalizer at LMC (the latter just to be sure) in that order
=> Restarted GEIRS readout, readout manager, scheduler and journalizer at LMC (in that order)
=> Initialized all units except MOS at instrument manager(again just to be sure)
=> Started readout manager: took image without saving OK, took image with saving, OK (luci2.20140204.0001)
=> Loaded Dark_RN_NoCU.xml into UI, unchecked the instrument tab on the UI
=> 8:10 UT Started Dark_RN_NoCU.xml
=> 10:31 Script finished, files luci2.20140204.0002-0057

February 19 2014

Tests by JH

Times are UT, files are luci2.20140219.iiii
07:30 Startup system
ROT=210:
FW1_repro: 10-19
FW2_repro: 20-29
Grat_select_repro:
08:25 Doesn't work, can only select blind mask nothing else???
=> Only blind mask was activated, did it, ini MOS+IM, stop/restart scheduler,journalizer => works
Grat_select_repro: 30-37
=> Dennis change of the calib-scripts with the CalUnit gives an error after the end of the script (when trying to move out the CU), script can also not be reloaded in scripting tool
09:05 Had to stop with tests, LBT reopened
09:50 continue, LBT closed again,
09:52 Screen frozen while editing...terminated nxclient and restart
Line width/dispersion:
G210z, 0.95mue, Argon, on/off, 10sec, #38/39
G210J, 1.25mue, Argon on/off, 10sec, #40/41
G210H, 1.265mue, Argon on/off, 10sec, #42/43
G210Ks, 2.17mue, Argon on/off, 10sec, #44/45
G200zJ, 0.95mue, Argon on/off, 10sec, #46/47
G200zJ, 1.12mue, Argon on/off, 10sec, #48/49
G200HK, 1.65mue, Argon on/off, 10sec, #50/51
G200HK, 2.17mue, Argon on/off, 10sec, #52/53
ROT=300:
Grat_select_repro: #54-62
FW1_repro: #63-72
FW2_repro: #73-82
Rot=120
FW1_repro: #83-92
FW2_repro: #93-96
FW2 got stuck after #96
Recover, OK
FW2_repro: #97-99
FW2 ggot stuck again, doesn"t like to go over from Ks to HKspec at that rot, need to modify script for tomorrow
Grat_select_repro: #100
same problem with FW2, need to modify script as well.
=> done for today, to be continued tomorrow if possible

February 20 2014

Test by DT

GEIRS gui came up fine, can take images from it.
Execution panel, LUCI2 and READ2 guis came up fine.
Cannot read detector from READ2 panel.
I get a popup that says "broken pipe" and a log message of:
"RemoteException occurred in server thread..."

LMC:
Stop 2 GEIRS Readout
Stop 2 Readout Manager
Stop Scheduler
Start 2 GEIRS Service (error starting process)
Start 2 Readout Manager (error starting process)
Start Scheduler (error starting process)

Can now read images from READ2 gui

Insert Calibration Unit
N3.75, mirror, P_gam (1.094um) + clear
Halo3 on to test exposure time...100 ADU in 2.75s
HeI (1.088um) to test...125 ADU in 2.75s
...back to P_gam

Data: luci2.20140220.NNNN.fits
All are LIR mode. Data taken in pairs for gain measurement.

These taken manually from READ2 GUI:
0001 0002 2.7399s
0003 0004 2.75
0005 0006 3.00
0007 0008 3.50
0009 0010 4.00
0011 0012 5.00
0013 0014 7.00
0015 0016 9.00
0017 0018 12.00
0019 0020 15.00
0021 0022 20.00
0023 0024 25.00
0025 0026 30.00
0027 0028 40.00
0029 0030 50.00

The interspersed 2.75s exposures are not really necessary.
Make a script...Gain_LIR.xml (in technical_scripts)
0031 0032 1.00 (2.7399)
0033 0034 2.75 ~100 ADU
0035 0036 3.0
0037 0038 3.5
0039 0040 4.0
0041 0042 4.5
0043 0044 5.0
0045 0046 6.0
0047 0048 7.0
0049 0050 8.0
0051 0052 10.0
0053 0054 12.0
0055 0056 15.0
0057 0058 20.0
0059 0060 25.0
0061 0062 30.0
0063 0064 40.0
0065 0066 50.0
0067 0068 70.0
0069 0070 90.0
0071 0072 110.0 ~4000 ADU
0073 0074 130.0

Make a script...Gain_MER.xml (in technical_scripts)
0075 0076 1.0 (6.8515)
0077 0078 7.0
0079 0080 9.0
0081 0082 12.0
0083 0084 16.0
0085 0086 20.0
0087 0088 26.0
0089 0090 34.0
0091 0092 45.0
0093 0094 58.0
0095 0096 76.0
0097 0098 100.0
0099 0100 130.0

Re-run Gain_LIR.xml
0101 0102 1.00 (2.7399)
0103 0104 2.75 ~100 ADU
0105 0106 3.0
0107 0108 3.5
0109 0110 4.0
0111 0112 4.5
0113 0114 5.0
0115 0116 6.0
0117 0118 7.0
0119 0120 8.0
0121 0122 10.0
0123 0124 12.0
0125 0126 15.0
0127 0128 20.0
0129 0130 25.0
0131 0132 30.0
0133 0134 40.0
0135 0136 50.0
0137 0138 70.0
0139 0140 90.0
0141 0142 110.0 ~4000 ADU
0143 0144 130.0

05:26 - End of DThompson calibrations.

Tests by JH

08:30 Winds high at LBT, JH calibs cont'd
Startup system without problems
Modified script for FW2-repro and grating-tilt repro to avaoid to go over H/K
ROT=30
FW1_repro: #145-154
FW2_repro:#155-164
"Error during setup verification:"coud not find grating name"
Grat_tilt_repro_G200: #165-168
Grat didn't want to tilt at 1.93mue at 5th exp., retry
Grat_tilt_repro_G200: #169-174
09:10 Had to stop, wind went down, they open the enclosure
09:45 Wind rose up again, cont'd
Grat_tilt_repro_G200: #175-180
Grat_select_repro: G210 will not stabilize at 1.25mue at 1st exp.,
retry, #181-186, G210/1.25mue not stab again, retry, failed again,
set lcen to 1.2mue in script, retry, OK
Grat_select_repro: #187-195
ROT=120
Grat_select_repro: #196-204
Grat_tilt_reproG210: #205-208, G210 will again not tilt, this
time at 2.17mue, retry, #209-214, OK
FW2_repro: #215-224
11:30 Done

-- %USERSIG{JochenHeidt - 2014-02-21}%

Comments

 
Topic revision: r3 - 03 May 2015, JochenHeidt
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback